A 20 g sample of ice at -10 °C is mixed with 100 g water at 80 °C. Calculate the

final temperature of the mixture.


1
Expert's answer
2021-11-04T00:31:55-0400

The specific heat capacity of ice (Cice) is 2.108 J g−1 °C−1

The specific heat capacity of water (Cwater) is 4.187 J g−1 °C−1

The heat of fusion of ice (ΔHf) is 333.55 J g−1

T1 = -10°C

T2 = 0°C

T3 = 80°C


Solution:

This problem can be summarized thusly:

1) Ice is heated from -10°C to 0°C.

The heat is calculated by using the specific heat of ice (Cice) and the equation: q = m × C × ΔT

q1 = mice × Cice × (T2 - T1)

q1 = (20 g) × (2.108 J g−1 °C−1) × (0 - (-10))°C = 421.6 J

q1 = 421.6 Joules


2) Ice is melted at 0°C.

The heat is calculated by multiplying the mass of ice by the heat of fusion of ice (ΔHf).

q2 = ΔHf × mice = (333.55 J g−1) × (20 g) = 6671 J

q2 = 6671 Joules


3) 20 g of water at 0°C is heated to the final temperature of the mixture.

The heat is calculated by using the specific heat of water (Cwater) and the equation: q = m × C × ΔT

q3 = mice × Cwater × (Tf - T2)

q3 = (20 g) × (4.187 J g−1 °C−1) × (Tf - 0)°C = 83.74 J × (Tf - 0)

q3 = 83.74 × Tf Joules


4) 100 g of water at 80°C is cooled to the final temperature of the mixture.

The heat is calculated by using the specific heat of water (Cwater) and the equation: q = m × C × ΔT

q4 = mwater × Cwater × (Tf - T3)

q4 = (100 g) × (4.187 J g−1 °C−1) × (Tf - 80)°C = 418.7 J × (Tf - 80)

q4 = 418.7 × Tf - 33496 Joules


Now just sum all of these and the sum must equal to zero; that is, heat gained by ice = heat lost by water.

q1 + q2 + q3 + q4 = 0

(421.6) + (6671) + (83.74 × Tf) + (418.7 × Tf - 33496) = 0

(-26403.4) + (502.44 × Tf) = 0

502.44 × Tf = 26403.4

Tf = 52.55°C


Answer: The final temperature of the mixture is 52.55°C